# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ Abstract
When making route decisions, travelers may engage in a certain degree of reasoning about what the others will do in the upcoming day, rendering yesterday’s shortest routes less attractive. This phenomenon was manifested in a recent virtual experiment that mimicked travelers’ repeated daily trip-making process. Unfortunately, prevailing day-to-day traffic dynamical models failed to faithfully reproduce the collected flow evolution data therein. To this end, we propose a day-to-day traffic behavior modeling framework based on the Cognitive Hierarchy theory, in which travelers with different levels of strategic-reasoning capabilities form their own beliefs about lower-step travelers’ capabilities when choosing their routes. Two widely-studied day-to-day models, the Network Tatonnement Process dynamic and the Logit dynamic, are extended into the framework and studied as examples. Calibration of the virtual experiment is performed using the extended Network Tatonnement Process dynamic, which fits the experimental data reasonably well. We show that the two extended dynamics have multiple equilibria, one of which is the classical user equilibrium. While analyzing global stability is intractable due to the presence of multiple equilibria, local stabilities near equilibria are developed analytically and verified by numerical experiments. General insights on how key parameters affect the stability of user equilibria are unveiled.
Keywords: day-to-day traffic dynamics; cognitive hierarchy; strategic thinking; route choice behavior; experiment calibration
@@ 1 Introduction
Wardrop’s first principle characterizes a steady-state User Equilibrium (UE) wherein no driver can reduce their travel time by unilaterally changing the route. However, due to the ever-changing nature of a transportation network (e.g., induced by traffic incidents, capacity modifications, and network structure changes), traffic states may always be in a disequilibrium state (Kumar and Peeta, 2015). To explain how travelers adjust their route-choice behaviors and predict those out-of-equilibrium states, various day-to-day traffic-network flow models have been proposed in the literature over the years.
@@
Most day-to-day models, either continuous or discrete, focus on modeling the impact of past flow evolution on travelers’ route choice behavior of today. The differences among the experienced travel times are the internal driving force that induces the flow change. For example, a rational behavior adjustment process (RBAP) proposed in Yang and Zhang (2009) generalizes a kind of flow evolution process along which the total travel cost of today will decrease based on the experienced cost of yesterday. A bunch of day-to-day models, including the proportional-switch adjustment (Smith, 1984), the projected dynamical system (Nagurney and Zhang, 1997), the network tatonnement process (Friesz et al., 1994), the evolutionary traffic dynamic (Sandholm, 2010) and the simplex gravity flow dynamic (Smith, 1983), can be categorized into this framework. Some studies also use the learning behavior (specifically an exponential smoothing filter) to model the impact of the whole historical cost trajectory on today’s route choice (Horowitz, 1984; Bie and Lo, 2010; Xiao et al., 2016, 2019).
@@
Largely ignored in the existing modeling efforts is the fact that travelers may predict what other travelers will do in the future (upcoming day). Substantial evidence in psychological studies demonstrates that when making decisions, people engage in a certain degree of reasoning about what others will do, i.e., a theory of mind, the ability to understand another person’s mental state (Lo, 2017). In a route-choice context, suppose now a transportation network is in a disequilibrium state, and travelers in the costlier route have the incentive to switch to the shortest route with the minimum travel time. A traveler may reason like: “if many travelers choose the shortest route, then I will try to avoid being on that route”. This phenomenon was somehow manifested in a recent day-to-day virtual experiment (Ye et al., 2018), but the models calibrated in that study did not take it into account. Without modeling such a logic, the prevailing day-to-day models are more like a way of calculating the final equilibrium (e.g., Hazelton and Watling, 2004) rather than explaining the inherent reasons for traffic flow evolution. Intuitively, if all the selfish travelers have yesterday’s complete information and do not consider what the others will do, the shortest one will be the only choice. Under this circumstance, the flow evolution trajectory may oscillate permanently among different routes.
@@
To model the travelers’ prediction behaviors, this paper employs the idea of Cognitive Hierarchy (CH) theory from Camerer et al. (2004), in which travelers form their beliefs of their opponents using an iterated reasoning process. In the CH theory, the heterogeneous players differ in their strategic-reasoning levels (i.e., cognitive capacities): while lower-step players do not carefully think through the whole game, the higher-step players would try to reap benefits by predicting how these lower-level “careless” players respond to the current situation. The CH model provided superior explanations in many game-theoretic settings, including guessing games (Costa-Gomes and Crawford, 2006) and extensive-form games (Ho and Su, 2013). It can also explain many phenomena in economics, marketing, and operation management, including market entry competition (Goldfarb and Xiao, 2011) and capacity allocation games (Cui and Zhang, 2017). It is worth mentioning that although the idea stemmed from Camerer et al. (2004), our proposed model’s novelty lies in its dynamic nature. In contrast, most applications of the CH model in the literature are static one-shot games such as the p-beauty game.
@@
In the literature, two pieces of work have recognized and modeled such prediction behavior; see He and Liu (2012) and He and Peeta (2016). However, they still have their limitations. The former work predicted the potential future congestion only once, and the prediction memory gradually vanished as time wore on. The latter work focused on the case where all the travelers are homogenous 1-step travelers; i.e., they all believe that the other travelers except herself are 0-step ones who myopically switch to shorter routes based on previous experience. Moreover, the model was developed using ordinary differential equations. Despite their tractability for mathematical analysis, the continuous models bring challenges in calibration with real data because the real-world system of repeated daily trips is discrete. Some properties (e.g., stability) in continuous time can only be considered a weaker result compared to the discrete-time counterpart (Li et al., 2018). Under some scenarios, conclusions drawn from continuous- and discrete-time systems may even contradict each other (Ye et al., 2021).
@@
In this paper, we establish a general framework of day-to-day network flow dynamics based on the CH theory to analyze travelers’ dynamic routing behaviors. In this framework, we allow the travelers to be heterogeneous in their strategic-reasoning capabilities and to form their own beliefs about other travelers’ strategic-reasoning capabilities and actions. We extend two widely-studied dynamics in the literature, the network tatonnement process (NTP) dynamic (for deterministic UE, DUE) and the Logit dynamic (for stochastic UE, SUE), into our CH framework and call them CH-NTP and CH-Logit dynamic, respectively. A previous online virtual experiment in Ye et al. (2018), which conventional day-to-day models could not fit, can now be calibrated by our proposed CH-NTP dynamic reasonably well. Moreover, the models predict that strategic-reasoning behaviors would lead the system to multiple equilibria, one of which being the classical user equilibrium. Jacobian matrices at the equilibrium points are derived and used to form analytical criteria on local stability around a user equilibrium. Theoretical results and numerical experiments unveiled several insights into how key parameters affect the local stability.
@@
The rest of the paper is organized as follows. The following section describes the general modeling framework. The CH-NTP dynamic is presented in Section 3 as the DUE example of the framework. Calibration of the real-world experiment with CH-NTP dynamic is furnished in Section 4. Section 5 provides numerical experiments regarding the CH-NTP dynamic. As the SUE example, the CH-Logit dynamic is presented in Section 6. Conclusions are drawn in Section 7 with discussions on future research directions.
@@ 2.2 Flow update rules
On day $t$, the route flow pattern of $k$-step is denoted as $\boldsymbol{x}^{k,(t)}$, and the aggregate observed route flow pattern $\tilde{\boldsymbol{x}}^{(t)}=\sum_{k}{\boldsymbol{x}^{k,(t)}}$. By observing $\tilde{\boldsymbol{x}}^{(t)}$ (or receiving it from the advanced traveler information system), a $k$-step traveler before trip will predict the route flow pattern on ($t+1$)-th day as per her cognitive capacity level (see subsequent Section 2.3). The predicted route flow pattern and its corresponding travel cost pattern are denoted as $\boldsymbol{\pi}^{k,(t+1)}$ and $\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)$, respectively. With $\boldsymbol{x}^{k,(t)}$ and $\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)$, the flow pattern of $k$-step travelers on the next day $t+1$ can be updated as:
@@
where $\boldsymbol{y}$ is the “target” flow pattern in $\Omega_{p^{k}}$ determined by a specific day-to-day dynamical model parameterized by $\zeta$; and the exponential-moving-average coefficient, $\alpha\in(0,1)$, reflecting travelers’ inertia or reluctance to change. (The larger the $\alpha$, the smaller the inertia.) Note that the predicted flow pattern for any $k$, $\boldsymbol{\pi}^{k,(t+1)}$, is always in $\Omega_{1}$.
Our model does not incorporate memory and learning processes in travelers’ decision-making. If we were to include memory effects, travelers’ route choices might be influenced by a weighted average of past experiences rather than just the most recent one (Horowitz, 1984; Bie and Lo, 2010; Xiao et al., 2016; Ye et al., 2021). By abstracting away memory and learning, we can more clearly isolate and analyze the impacts of travelers’ strategic thinking behavior on network dynamics. This simplification allows for greater analytical tractability and easier interpretation of results.
@@
By describing the dynamic in such a general way, most day-to-day models in the literature can be applied to the framework. For example, the day-to-day models satisfying the RBAP property (Yang and Zhang, 2009) can be applied, i.e., any $\boldsymbol{y}[\cdot]$ satisfying:
$\boldsymbol{y}\left[\boldsymbol{x}^{k,(t)},\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right);\zeta\right]\left\{\begin{aligned} &\in\Psi^{k,(t)},\Psi^{k,(t)}\neq\O\\ &=\boldsymbol{x}^{k,(t)},\Psi^{k,(t)}=\O\end{aligned}\right.,$ (2)
where $\Psi^{k,(t)}\equiv\left\{{\boldsymbol{y}|\boldsymbol{y}\in\Omega_{p^{k}},\boldsymbol{y}^{T}\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)<(\boldsymbol{x}^{k,(t)})^{T}\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)}\right\}$ is the feasible route flow set of $k$-step for which the $k$-step travelers’ total travel costs based on the predicted costs decreases.
@@
Moreover, applications of the framework are not confined to deterministic day-to-day models; i.e., the $\boldsymbol{y}[\cdot]$ can also be in a stochastic sense. For example, Logit-based stochastic day-to-day models contained in the “SUE version” of the RBAP framework (Xiao et al., 2019) can be adopted. In this paper, we study two typical examples for the DUE and SUE, respectively.
@@ 2.3 Cognitive hierarchy levels
It is clear that if the network only contains the 0-step travelers, it degenerates into a conventional day-to-day model. (In a CH model for the one-shot games, 0-step players were assumed to either randomize equally across all strategies or choose a salient strategy using ex-ante information (Camerer et al., 2004). Here we adopt the latter idea.)
As per the CH theory (Camerer et al., 2004), the $k(k\geq 1)$-step travelers try to take advantage by predicting how lower-step players respond to the current flow pattern, but are overconfident and do not realize others are using exactly as many thinking steps as they are. Denote $q_{k}^{h}=\frac{p^{h}}{\sum_{i=0}^{k-1}{p^{i}}}$ as $k$-step travelers’ belief about the normalized proportion of $h$-step travelers and $q_{k}^{h}=0$ for $\forall h\geq k$. As consistent with the CH theory, such a setting means that travelers do not know the exact distribution of lower-step travelers — they only confidently assume the normalized distribution. (This setting was also adopted in other studies; see, e.g., Cui and Zhang (2017) that apply the CH theory to the capacity allocation games in the operation management field.)
@@
where $\hat{\alpha}\in(0,1)$ and $\hat{\zeta}$ are the predicted coefficients in the travelers’ minds.
@@
Based on the predicted flow pattern $\boldsymbol{\pi}^{2,(t+1)}$, the 2-step travelers update their route choices as follows:
@@
Substantial experiments in the behavioral game theory found that the average thinking step across all the players lies between 1 and 2 (Camerer et al., 2004; Chong et al., 2016). Hence, the rest of this paper focuses on the case where $|K|\leq 3$. Note that those experiments were designed for one-shot games with a relatively small number of subjects, in contrast to the dynamic setting with many subjects in this paper. It is also worth mentioning that the methodology in this paper can also be applied to cases where $|K|>3$.
2.4 Model interpretation and practical relevance
2.4.1 Interpretations on the strategic thinking
Strategic thinking behavior is the core component of our proposed modeling framework. As stemmed from the CH theory (Camerer et al., 2004), our model assumes travelers know other travelers’ strategic levels (represented by $q_{k}^{h}$). This assumption may seem unrealistic at first glance since individual travelers are unlikely to have explicit information about the distribution of strategic sophistication among other road users. However, it can be interpreted on several grounds:
  1. 1.
    Assuming knowing lower-step travelers’ strategic levels serves as a simplified heuristic that represents the underlying cognitive processes travelers use to anticipate others’ behavior. While individual travelers may not explicitly possess this information, our model captures the essence of strategic thinking in aggregate. This approach is validated by the model’s improved fit to observed data (see Section 4), suggesting it effectively captures important aspects of real-world strategic behavior in transportation networks. Notably, even the 1-step model ($|K|=2$) provides considerable improvements while requiring only one additional parameter compared to traditional day-to-day models. The fact that a multi-class day-to-day model with additional parameters but without strategic prediction failed to reproduce the experimental data further supports that our model’s success stems from its ability to capture strategic thinking, not merely from increased parameter flexibility.
  2. 2.
    The $q_{k}^{h}$ values might be interpreted as aggregate beliefs held by a class of travelers, rather than exact knowledge possessed by individuals. This aggregate approach allows us to model the overall effect of strategic thinking without making strong claims about individual cognition. The aggregate beliefs represented by $q_{k}^{h}$ can be seen as emergent properties of the system, arising from the collective experience of travelers over time, rather than explicit individual knowledge.
  3. 3.
    Our modeling framework aligns with the “theory of mind” concept from cognitive psychology - the ability to attribute mental states and intentions to others (Lo, 2017). In our transportation context, the strategic levels and associated beliefs represent simplified versions of travelers’ theory of mind about other road users.

### v2

@@ Abstract
When making route decisions, travelers may engage in a certain degree of reasoning about what the others will do in the upcoming day, rendering yesterday’s shortest routes less attractive. This phenomenon was manifested in a recent virtual experiment that mimicked travelers’ repeated daily trip-making process. Unfortunately, prevailing day-to-day traffic dynamic models failed to faithfully reproduce the collected flow evolution data therein. To this end, we propose a day-to-day traffic behavior modeling framework based on the Cognitive Hierarchy theory, in which travelers with different levels of strategic-reasoning capabilities form their own beliefs about lower-step travelers’ capabilities when choosing their routes. Two widely-studied day-to-day models, the Network Tatonnement Process dynamic and the Logit dynamic, are extended into the framework and studied as examples. Calibration of the virtual experiment is performed using the extended Network Tatonnement Process dynamic, which fits the experimental data reasonably well. Our analysis reveals that both extended dynamics exhibit multiple equilibria, one of which corresponds to the classical user equilibrium. We further analyze and characterize these non-user-equilibrium states. While analyzing global stability is intractable due to the presence of multiple equilibria, local stability criteria near equilibria are developed analytically. General insights on how key parameters affect the stability of user equilibria are unveiled.
Keywords: day-to-day traffic dynamics; cognitive hierarchy; strategic thinking; route choice behavior; multiple equilibria; experiment calibration
@@ 1 Introduction
Wardrop’s first principle characterizes a steady-state User Equilibrium (UE) wherein no driver can reduce their travel time by unilaterally changing the route. However, due to the ever-changing nature of a transportation network (e.g., induced by traffic incidents, capacity modifications, and network structure changes), traffic states may always be in a disequilibrium state (Kumar and Peeta 2015). To explain how travelers adjust their route-choice behaviors and predict those out-of-equilibrium states, various day-to-day traffic-network flow models have been proposed in the literature over the years.
@@
Most day-to-day models, whether continuous or discrete, concentrate on modeling how the past evolution of traffic flow influences travelers’ route choice behavior today. The differences among the experienced travel times are the internal driving force that induces the flow change. For example, a rational behavior adjustment process (RBAP) introduced by Yang and Zhang (2009) generalizes a flow evolution process wherein the total travel cost today will decrease based on yesterday’s experienced cost. This framework encompasses numerous day-to-day models, including the proportional-switch adjustment (Smith 1984), the projected dynamical system (Nagurney and Zhang 1997), the network tatonnement process (Friesz et al. 1994), the evolutionary traffic dynamic (Sandholm 2010), and the simplex gravity flow dynamic (Smith 1983). Some studies model travelers’ memory processes by updating the perceived cost through an exponential moving average of past perceived and actual costs (Horowitz 1984, Cantarella and Cascetta 1995, Bie and Lo 2010, Xiao, Yang, and Ye 2016, Ye, Xiao, and Yang 2021).
@@
Largely ignored in the existing modeling efforts is the fact that travelers may predict what other travelers will do in the future (upcoming day). Substantial evidence in psychological studies demonstrates that when making decisions, people engage in a certain degree of reasoning about what others will do. This reasoning process is referred to as “theory of mind,” the ability to understand another person’s mental state (Lo 2017). In a route-choice context, suppose now a transportation network is in a disequilibrium state, and travelers in the costlier route have the incentive to switch to the shortest route with the minimum travel time. A traveler may reason like: “if many travelers choose the shortest route, then I will try to avoid being on that route”. This phenomenon was somehow manifested in a recent day-to-day virtual experiment (Ye, Xiao, and Yang 2018), but the models calibrated in that study did not take it into account. Intuitively, if all travelers acted with perfect rationality, using only yesterday’s complete information without predicting what others might do today, they would all choose yesterday’s shortest route. Such behavior would result in permanent flow oscillations between routes. Yet real-world traffic patterns rarely exhibit such extreme fluctuations, suggesting other behavioral factors at play. Previous research has addressed this discrepancy through modeling travelers’ inertia — their reluctance to change routes completely from day to day. However, the strategic prediction of others’ route choices remains largely unexplored.
@@
To model the travelers’ prediction behaviors, this paper employs the idea of Cognitive Hierarchy (CH) theory from Camerer, Ho, and Chong (2004), in which travelers form their beliefs of their opponents using an iterated reasoning process. In the CH theory, the heterogeneous players differ in their strategic-reasoning levels (i.e., cognitive capacities). Lower-step players do not carefully think through the whole game. Meanwhile, higher-step players would try to reap benefits by predicting how these lower-level “careless” players respond to the current situation. The CH model provided superior explanations in many game-theoretic settings, including guessing games (Costa-Gomes and Crawford 2006) and extensive-form games (Ho and Su 2013). It can also explain many phenomena in economics, marketing, and operation management, including market entry competition (Goldfarb and Xiao 2011), platform competition (Hossain and Morgan 2013), information disclosure in movie markets (Brown, Camerer, and Lovallo 2013) and capacity allocation games (Cui and Zhang 2017). It is worth mentioning that although the idea stemmed from Camerer, Ho, and Chong (2004), our proposed model’s novelty lies in its dynamic nature. In contrast, most applications of the CH model in the literature are static one-shot games such as the $p$-beauty game.
@@
In the literature, two pieces of work have modeled such prediction behavior; see He and Liu (2012) and He and Peeta (2016). However, both have their limitations. The former work predicted future congestion only once, with prediction memory gradually vanishing over time. The latter work focused on the case where all the travelers are homogenous 1-step travelers who believe others are 0-step travelers making myopic switches based on previous experience. Moreover, this model uses ordinary differential equations which, despite their tractability for mathematical analysis, bring challenges in calibrating with real-world discrete daily trip data. Stability in continuous time may represent weaker results compared to the discrete-time counterpart (Li, Liu, and Nie 2018). Under certain scenarios, conclusions from continuous- and discrete-time systems may even contradict each other (Ye, Xiao, and Yang 2021).
@@
In this paper, we establish a general framework of day-to-day network flow dynamics based on the CH theory to analyze travelers’ repeated routing behaviors. In this framework, travelers are heterogeneous in their strategic-reasoning capabilities and form their own beliefs about other travelers’ strategic-reasoning capabilities and actions. We extend two widely-studied dynamics in the literature into our CH framework: the Network Tatonnement Process (NTP) dynamic for deterministic UE (DUE) and the Logit dynamic for stochastic UE (SUE). The two extended models are referred to as CH-NTP and CH-Logit dynamics, respectively. A previous online virtual experiment in Ye, Xiao, and Yang (2018), which conventional day-to-day models could not fit, can now be calibrated by our proposed CH-NTP dynamic reasonably well. Moreover, the two models predict that strategic-reasoning behaviors would lead the system to multiple equilibria, one of which being the classical UE. We formally characterize these equilibria and show that those which differ from UE emerge when travelers excessively anticipate others’ behaviors. Jacobian matrices at the equilibrium points are derived and used to form analytical criteria on local stability around UE. Theoretical results and numerical experiments unveiled several insights into how key parameters affect the local stability.
@@
The rest of the paper is organized as follows. The following section describes the general modeling framework. The CH-NTP dynamic is presented in Section 3 as the DUE example of the framework. Calibration of the real-world experiment with CH-NTP dynamic is furnished in Section 4. As the SUE example, the CH-Logit dynamic is presented in Section 6. Conclusions are drawn in Section 7 with discussions on future research directions.
@@ 2.2 Flow update rules
On day $t$, the route flow pattern of $k$-step is denoted as $\boldsymbol{x}^{k,(t)}$, and the aggregate observed route flow pattern $\tilde{\boldsymbol{x}}^{(t)}=\sum_{k}{\boldsymbol{x}^{k,(t)}}$. Throughout the paper, we use the notation $\sim$” above a variable to denote the aggregation of route flows across all classes, resulting in a total flow pattern. By observing $\tilde{\boldsymbol{x}}^{(t)}$ (or receiving it from the advanced traveler information system), a $k$-step traveler before trip will predict the route flow pattern on ($t+1$)-th day as per her cognitive capacity level (see subsequent Section 2.3). The predicted route flow pattern and its corresponding travel cost pattern are denoted as $\boldsymbol{\pi}^{k,(t+1)}$ and $\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)$, respectively. With $\boldsymbol{x}^{k,(t)}$ and $\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)$, the flow pattern of $k$-step travelers on the next day $t+1$ can be updated as:
@@
where $\boldsymbol{y}$ represents the “target” flow pattern in $\Omega_{p^{k}}$ determined by a specific day-to-day dynamical model parameterized by $\zeta$, and $\alpha\in(0,1]$ is a parameter that reflects travelers’ inertia or reluctance to change. Note that the predicted flow pattern for any $k$, $\boldsymbol{\pi}^{k,(t+1)}$, is always in $\Omega_{1}$.
Eq. (1) describes a general day-to-day adjustment process in which travelers partially adjust their route choices toward a target flow pattern $\boldsymbol{y}$ based on current predicted costs, with $\alpha$ controlling the extent of this adjustment (larger $\alpha$ indicates smaller inertia). This convex combination structure provides a memory and learning process by weighing current decisions against previous ones, a form widely adopted in day-to-day models (Friesz et al. 1994, He, Guo, and Liu 2010, Han and Du 2012, Guo et al. 2015) to represent the learning and adapting process before reaching a steady state (Wu, Yin, and Lynch 2024). Through instantiating different target flow patterns, we can derive many traditional day-to-day models in the literature from this general formulation.
While Eq. (1) provides a standard form of memory and learning process, we deliberately omit the more complex memory effects presented in previous day-to-day models (Horowitz 1984, Cantarella and Cascetta 1995, Bie and Lo 2010, Xiao, Yang, and Ye 2016, Ye, Xiao, and Yang 2021). These studies represent another line of research that explicitly models the updates of travelers’ perceptions of travel costs and the resulting route choices (Wu, Yin, and Lynch 2024). Our decision to implement the memory and learning mechanism in Eq. (1) rather than the more complex memory process is based on the following considerations: (i) it aligns with our empirical findings from experimental data (Ye, Xiao, and Yang 2018) analyzed in Section 4 where incorporating the complex memory process showed minimal improvements in explanatory power; (ii) it helps distinguish the effects of strategic prediction from those of more complicated memory effect; and (iii) it maintains analytical tractability while still capturing important behavioral insights, including the emergence of multiple equilibria (Section 3) and substantially improved goodness-of-fit to experimental data (Section 4.2).
@@
By describing the dynamic in such a general way, most day-to-day models in the literature can be applied to the framework. For example, the day-to-day models satisfying the RBAP property (Yang and Zhang 2009) can be applied, i.e., any $\boldsymbol{y}[\cdot]$ satisfying:
$\boldsymbol{y}\left[\boldsymbol{x}^{k,(t)},\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right);\zeta\right]\left\{\begin{aligned} &\in\Psi^{k,(t)},\Psi^{k,(t)}\neq\emptyset\\ &=\boldsymbol{x}^{k,(t)},\Psi^{k,(t)}=\emptyset\end{aligned}\right.,$ (2)
where $\Psi^{k,(t)}\equiv\left\{{\boldsymbol{y}|\boldsymbol{y}\in\Omega_{p^{k}},\boldsymbol{y}^{T}\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)<(\boldsymbol{x}^{k,(t)})^{T}\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)}\right\}$ is the feasible route flow set of $k$-step for which the $k$-step travelers’ total travel costs based on the predicted costs decreases. This dynamic process describes how, at each time step $t$, the current path flow pattern of $k$-step travelers, $\boldsymbol{x}^{k,(t)}$, evolves toward a target flow pattern $\boldsymbol{y}$ that reduces the total travel cost of the $k$-step travelers based on their predicted cost $\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)$. If such a cost-reducing flow $\boldsymbol{y}$ does not exist (i.e., $\Psi^{k,(t)}$ is empty), the flow remains unchanged. This formulation defines the evolution direction from $\boldsymbol{x}^{k,(t)}$ as a set rather than a single point, allowing it to encompass many day-to-day models as special cases.
@@
Moreover, applications of the framework are not confined to deterministic day-to-day models; i.e., the $\boldsymbol{y}[\cdot]$ can also be in a stochastic sense. For example, Logit-based stochastic day-to-day models contained in the “SUE version” of the RBAP framework (Xiao et al. 2019) can be adopted. In this paper, we study two typical examples for the DUE and SUE, respectively.
@@ 2.3 Cognitive hierarchy levels
Here, $\tilde{\boldsymbol{x}}^{(t)}$ represents the aggregate observed route flow pattern on day $t$, which is the sum of route flows across all the classes. In a CH model for the one-shot games, 0-step players were assumed to either randomize equally across all strategies or choose a salient strategy using ex-ante information (Camerer, Ho, and Chong 2004). Here, we adopt the latter idea by modeling 0-step travelers using conventional day-to-day models without prediction. This approach offers two advantages: it incorporates conventional day-to-day models as special cases (when $p^{0}=1$) and provides a tractable, interpretable baseline for higher-level travelers to build their predictions upon.
As per the CH theory (Camerer, Ho, and Chong 2004), the $k(k\geq 1)$-step travelers try to take advantage by predicting how lower-step players respond to the current flow pattern. However, they are overconfident and do not realize others are using exactly as many thinking steps as they are. Denote $q_{k}^{h}=\frac{p^{h}}{\sum_{i=0}^{k-1}{p^{i}}}$ as $k$-step travelers’ belief about the normalized proportion of $h$-step travelers and $q_{k}^{h}=0$ for $\forall h\geq k$. As consistent with the CH theory, such a setting means that travelers do not know the exact distribution of lower-step travelers — they only confidently assume the normalized distribution. For practical implementation, modelers can treat these strategic level proportions ($p^{k}$) as parameters to be calibrated alongside other model parameters. The calibration can be performed by minimizing the root mean squared error (RMSE) between predicted and observed flow patterns across multiple days. We demonstrate this approach in Section 4.2.
@@
where $\hat{\alpha}\in(0,1)$ and $\hat{\zeta}$ are the predicted coefficients in travelers’ minds.
@@
Based on $\boldsymbol{\pi}^{2,(t+1)}$, the 2-step travelers update their route choices as follows:
@@
Substantial experiments in the behavioral game theory found that the average thinking step across all the players lies between 1 and 2 (Camerer, Ho, and Chong 2004, Chong, Ho, and Camerer 2016). Hence, the rest of this paper focuses on the case where $|K|\leq 3$. Note that these experiments were designed for one-shot games with few subjects (atomic players), unlike our dynamic setting with many non-atomic participants. It is also worth mentioning that the methodology in this paper can also be applied to cases where $|K|>3$.
2.4 Justifications and interpretations of applying the strategic thinking into the day-to-day traffic dynamics
Strategic thinking behavior is the core component of our proposed modeling framework. Stemming from the CH theory (Camerer, Ho, and Chong 2004), we model heterogeneous travelers with different strategic-reasoning capabilities. Higher-level travelers predict the actions of lower-level travelers but assume no one else possesses their level of sophistication. This fundamental premise has been successfully applied in various domains such as platform competition markets (Hossain and Morgan 2013), information disclosure in movie markets (Brown, Camerer, and Lovallo 2013), and capacity allocation games in operations management (Cui and Zhang 2017). We preserve this essential CH mechanism while adapting it to the non-atomic, dynamic day-to-day traffic setting.
2.4.1 Justifications for strategic thinking
We offer four reasons for incorporating strategic thinking into day-to-day traffic dynamics:
1. Our virtual experiment (which we will discuss in detail in Section 4) revealed an interesting phenomenon: during the early stages of the experiment, participants sometimes chose yesterday’s second-shortest route over the shortest route. This phenomenon contradicts conventional day-to-day models, which would predict travelers gravitating toward the shortest path. Instead, this phenomenon suggests strategic thinking: participants anticipated that others would choose yesterday’s shortest route and deliberately avoided it to escape potential congestion. Moreover, interviews with a subset of the 268 participants revealed that more than one-third described using such strategic route evaluations in their decision-making process.
2. Regarding model calibration for our experiment, we conducted ablation studies that removed higher-level strategic thinking ($k\geq 1$) and treated all travelers as 0-step players (as in conventional day-to-day models). The calibrated results consistently failed to reproduce the experimental data. This failure is shared by all the five conventional day-to-day models examined in Ye, Xiao, and Yang (2018). Such consistent failure across multiple models demonstrates the necessity of incorporating strategic thinking.
3. In addition to our experiment, Selten et al. (2007)’s comprehensive 200-period route choice experiments also demonstrated that participants explicitly engage in strategic prediction of others’ choices. The authors identified a “contrary” response mode where participants switch routes after experiencing good outcomes because they anticipated that others would be attracted to those routes in subsequent periods. Similar prediction behavior was also clearly documented in a more recent experiment conducted by Liu et al. (2020).
4. Beyond laboratory experiments, He and Liu (2012) analyzed real-world citywide detector data following the I-35W Bridge collapse in Minneapolis, Minnesota. The authors found that drivers made route choices based on predicted traffic flow patterns rather than merely reacting to experienced congestion. Notably, even immediately after the disruption, when drivers had not yet experienced the new traffic conditions, they already adjusted their routes by anticipating potential congestion on what would typically be considered shortest paths. This real-world evidence demonstrates that strategic thinking in route choice extends beyond laboratory settings to actual travel behavior.
2.4.2 Model interpretations
Following the CH theory, travelers form beliefs about the distribution of lower-level types using normalized proportions $q_{k}^{h}$ rather than exact population distributions. We interpret the normalized proportions as a simplified heuristic representing the “strength” of underlying cognitive processes travelers use to anticipate others’ behavior. These normalized proportions captures whether prediction tendencies are relatively stronger or weaker within the population. They might be viewed as aggregate beliefs held by a class of travelers, rather than exact knowledge possessed by individuals. Furthermore, our day-to-day model operates at a macroscopic level with non-atomic users. At this aggregate level, the proportions can be viewed as emergent properties arising from collective behavior. Finally, our modeling framework aligns with the “theory of mind” concept from cognitive psychology — the ability to attribute mental states and intentions to others (Lo 2017). In our transportation context, the strategic levels and associated beliefs represent a manifestation of travelers’ theory of mind about other road users.